# Distributionally Robust System Level Synthesis With Output Feedback Affine Control Policy

### v1

@@ Abstract
This paper studies the finite-horizon robust optimal control of linear systems subject to model mismatch and additive stochastic disturbances. Utilizing the system level synthesis (SLS) parameterization, we propose a novel SLS design using output-feedback affine control policy and extend it to a distributionally robust setting to improve system resilience by minimizing the cost function while ensuring constraint satisfaction against the worst-case uncertainty distribution. The scopes of model mismatch and stochastic disturbances are quantified using the 1-norm and a Wasserstein metric-based ambiguity set, respectively. For the closed-loop dynamics, we analyze the distributional shift between the predicted output-input response – computed using nominal parameters and empirical disturbance samples – and the actual closed-loop distribution, highlighting its dependence on model mismatch and SLS parameterization. Assuming convex and Lipschitz continuous cost functions and constraints, we derive a tractable reformulation of the distributionally robust SLS (DR-SLS) problem by leveraging tools from robust control and distributionally robust optimization (DRO). Numerical experiments validate the performance and robustness of the proposed approach.
@@ 1 Introduction
To enhance robustness and reduce conservatism in optimal control design, closed-loop feedback policies are generally preferred over open-loop control actions in both robust optimal decision-making and Model Predictive Control (MPC) frameworks [4, 17]. However, incorporating feedback policies directly into online MPC optimization introduces significant computational challenges, as it results in nonconvex optimization problems–even when the system dynamics and control laws are linear. For linear systems, other than the conventional MPC design focusing on optimizing controllers, the system level synthesis (SLS) framework provides a parameterization of the closed-loop system and alternatively focuses on directly optimizing the whole closed-loop performance by solving convex optimization problems and reconstructing control policies from the optimized closed-loop response via linear transformations [18]. This shift in perspective has facilitated progress in distributed and robust predictive control design [18, 19, 20].
@@
Motivated by the above discussions, this paper investigates SLS design with output feedback affine control policies for uncertain linear systems subject to model mismatch and additive disturbances. This setting presents several major challenges: 1) the use of affine output feedback policies complicates the construction of an SLS parameterization while maintaining computational tractability; 2) model mismatch and stochastic disturbances lead to inaccurate predictions of system dynamics, increasing the difficulty of designing effective feedback policies; 3) model mismatch induces distributional shifts between the true closed-loop behavior and the predicted closed-loop response, further complicating the solution of the DR control formulation.
To address the above challenges, we propose a novel DR-SLS method with output feedback affine policies. The contributions of this work are summarized as follows:
@@
  • A computationally tractable relaxation of the proposed DR-SLS problem is developed by incorporating tools from robust SLS and DRO, resulting a convex reformulation. Simulation results are provided to validate the effectiveness of the proposed approach.
@@ 2 Problem Formulation
where $x_{t}\in\mathbb{R}^{n}$ is the system state, $u_{t}\in\mathbb{R}^{m}$ is the control input, and $e_{t}\in\mathbb{R}^{q}$ is the innovation process. The matrices $A$, $B$, $L$, $C$ and $D$ are the respective system, input, Kalman, output and feedforward matrices, respectively. More details about the innovation form of linear systems can be found in [26, 27, 28].
@@
where $T$ is the length of prediction horizon, $h_{t}(\cdot)$ and $g_{t}(\cdot)$ are stage cost and constraint functions, respectively, $\pi(\cdot)$ is the control policy. In our design, $h_{t}$ and $g_{t}$ are assumed to be convex and Lipschitz continuous functions.
@@
In the upcoming sections, the SLS design for system (1) under control law (3) will be investigated. It is important to note that incorporating the affine control policy introduces computational challenges for existing SLS methods. Specifically, the presence of the bias term in the affine policy renders that a direct extension of the existing SLS methods developed for linear control policies, see [22, 23], becomes nonconvex and thus computationally demanding. In addition to developing an SLS design for system (1) under the assumption of perfect knowledge of system models, we further investigate the impact of model uncertainties for the proposed SLS design. To address the influence of model uncertainties, we propose a novel DR-SLS framework to improve the robustness of the optimal control solution against model mismatch and stochastic disturbances.
@@ 3.1 System Level Synthesis With Exact Model
and for any $\tau\in\mathbb{Z}_{+}$ (also called past horizon), it can be derived that
@@
Supposing that (ALC)(A-LC) is Schur stable and denoting $\lambda_{\max}$ ($|\lambda_{\max}|<1$) as the eigenvalue of (ALC)(A-LC) with the largest absolute value, it can be derived that
@@
$\mathbf{G}=\mathbf{CT}_{u}+\mathbf{D},\ \mathbf{y}_{0}=\mathbf{C}\Gamma\Psi\begin{bmatrix}\mathbf{u}_{0}^{-}\\ \mathbf{y}_{0}^{-}\end{bmatrix},\ \Theta=\mathbf{CT}_{e}+I.$ (10)
@@
By substituting $\mathbf{K}=\Phi_{u}\Phi_{y}^{-1}$, we obtain
@@
where the second equality exploits the property $\mathbf{G}\Phi_{u}=\Phi_{y}-I$ in (14).
Similarly, one has
@@
Considering the equality (19) and substituting $\mathbf{K}=\Phi_{u}\Phi_{y}^{-1}$ and $\mathbf{p}=\phi_{u}-\mathbf{K}\phi_{y}$ into the first equality in (18b) gives
$(I-\mathbf{GK})^{-1}\mathbf{Gp}=$ $\Phi_{y}\mathbf{G}(\phi_{u}-\Phi_{u}\Phi_{y}^{-1}\phi_{y})$ (21a)
$=$ $\Phi_{y}(\phi_{y}-(\Phi_{y}-I)\Phi_{y}^{-1}\phi_{y})=\phi_{y},$ (21b)
where the second equality builds upon the property $\mathbf{G}\phi_{u}=\phi_{y}$ and $\mathbf{G}\Phi_{u}=\Phi_{y}-I$ in (14).
@@
$(I-\mathbf{KG})^{-1}\mathbf{p}=$ $(I+\mathbf{K}(I-\mathbf{GK})^{-1}\mathbf{G})(\phi_{u}-\mathbf{K}\phi_{y})$
$=$ $(I+\Phi_{u}\mathbf{G})(\phi_{u}-\Phi_{u}\Phi_{y}^{-1}\phi_{y})$
$=$ $(\phi_{u}-\Phi_{u}\Phi_{y}^{-1}\phi_{y})+\Phi_{u}(\mathbf{G}\phi_{u}-\mathbf{G}\Phi_{u}\Phi_{y}^{-1}\phi_{y})$
$=$ $(\phi_{u}-\Phi_{u}\Phi_{y}^{-1}\phi_{y})+\Phi_{u}(\phi_{y}-(\Phi_{y}-I)\Phi_{y}^{-1}\phi_{y})$
$=$ $\phi_{u},$
where the first equality exploits the matrix inverse lemma (Sherman-Morrison-Woodbury formula) [30]. This completes the proof. $\square$
@@
where $\mathbb{P}^{\mathcal{M}}$ denotes the probability distribution of the closed-loop output-input response $\bm{\eta}$ with the system model $\mathcal{M}$, which influences the parameters $(\mathbf{G},\Theta,\mathbf{y}_{0})$, the functions $h(\cdot)$ and $g(\cdot)$ are constructed based on $h_{t}(\cdot)$ and $g_{t}(\cdot)$ defined in (2), respectively. The stochastic uncertainty considered in the above formulation is the innovation process vector $\mathbf{e}$.
@@ 3.2 SLS Parameterization With Model Mismatch
The SLS formulation proposed in Section 3.A is based on the assumption that the system model, i.e., the parameters $(\mathbf{G},\Theta,\mathbf{y}_{0},\mathbf{e})$, is perfectly known. However, obtaining the exact model information is nontrivial and challenging. It is more practical to consider that only an approximate model is available for SLS design.
In this section, the SLS formulation proposed in Theorem 1 of Section 3.A is extended to the case where only approximated parameters $(\hat{\mathbf{G}},\hat{\Theta},\hat{\mathbf{y}}_{0})$ are available. For ease of notation, the approximation error of $\mathbf{G},\Theta,\mathbf{y}_{0}$ are denoted as $\Delta:=\mathbf{G}-\hat{\mathbf{G}}$, $\tilde{\Theta}:=\Theta-\hat{\Theta}$, and $\tilde{\mathbf{y}}_{0}:=\mathbf{y}_{0}-\hat{\mathbf{y}}_{0}$, respectively. It is worth noting that among the above-listed parameters, $\mathbf{G}$ plays the most important role in influencing the SLS design since it will directly influence the SLS design via (14).
@@
Proof: Since $\hat{\Phi}_{y}$ and $\hat{\Phi}_{u}$ satisfy (14), it can be derived that
$\begin{bmatrix}I&-\mathbf{G}\end{bmatrix}\begin{bmatrix}\hat{\Phi}_{y}\\ \hat{\Phi}_{u}\end{bmatrix}=I-\Delta\hat{\Phi}_{u},$
@@
$\begin{bmatrix}I&-\mathbf{G}\end{bmatrix}\begin{bmatrix}\hat{\Phi}_{y}\\ \hat{\Phi}_{u}\end{bmatrix}\left(I-\Delta\hat{\Phi}_{u}\right)^{-1}=I.$ (26)
The above equality provides an option for selecting $\Phi_{y}=\hat{\Phi}_{y}(I-\Delta\hat{\Phi}_{u})^{-1}$ and $\Phi_{u}=\hat{\Phi}_{u}(I-\Delta\hat{\Phi}_{u})^{-1}$ that ensures $\mathbf{K}=\Phi_{u}\Phi_{y}^{-1}=\hat{\Phi}_{u}\hat{\Phi}_{y}^{-1}$.
Similarly, since $\hat{\phi}_{y}$ and $\hat{\phi}_{u}$ satisfy
@@
where the relationship in (29b) is built upon the fact that the control policy parameter $\mathbf{p}$ is identical for both cases of the true closed-loop parameterization and nominal closed-loop parameterization.
Denoting $r_{y}:=\phi_{y}-\hat{\phi}_{y}$ and $r_{u}:=\phi_{u}-\hat{\phi}_{u}$, and combining (28) and (29) result in
@@ 4.1 Distributionally Robust SLS Formulation
Solving the SLS-based problem (23) is generally intractable when assuming the existence of model mismatch $\{\Delta,\tilde{\Theta},\tilde{\mathbf{y}}_{0}\}$ and uncertain probability distribution $\mathbb{P}_{\mathbf{e}}$. One possible solution for solving (23) is adopting the sample average approximation (SAA) and certainty equivalence (CE) principles, in which the unknown system parameters $(\mathbf{G},\Theta,\mathbf{y}_{0})$ are replaced by their nominal values $(\hat{\mathbf{G}},\hat{\Theta},\hat{\mathbf{y}}_{0})$ and the true innovation process distribution $\mathbb{P}_{\mathbf{e}}$ is approximated by its empirical counterpart $\overline{\mathbb{P}}_{\mathbf{e}}$. In this way, the true closed-loop response distribution $\mathbb{P}^{\mathcal{M}}$ is replaced by its empirical predictive distribution $\overline{\mathbb{P}}^{\hat{\mathcal{M}}}$.
@@
In the following, a distributionally robust SLS (DR-SLS) design will be presented, considering the uncertainties introduced by the model mismatch between $(\mathbf{G},\Theta,\mathbf{y}_{0})$ and $(\hat{\mathbf{G}},\hat{\Theta},\hat{\mathbf{y}}_{0})$ as well as the estimation error of the empirical innovation process distribution $\overline{\mathbb{P}}_{\mathbf{e}}$. Before presenting the main results of this section, the following definitions are introduced.
@@
Given $N$ samples of innovation process $\{\mathbf{e}_{i}\}_{i=1}^{N}$, the empirical distribution $\overline{\mathbb{P}}_{\mathbf{e}}$ can be constructed as $\overline{\mathbb{P}}_{\mathbf{e}}:=\frac{1}{N}\sum_{i=1}^{N}\delta(\mathbf{e}_{i})$, where $\delta(\cdot)$ is the Dirac distribution. Accordingly, the empirical predictive distribution of the output-input response $\bm{\eta}$ with the nominal model $\mathcal$$\hat${M}} is

### v2

@@ Abstract
This paper studies the finite-horizon robust optimal control of constrained linear systems subject to model mismatch and additive stochastic disturbances. Utilizing the system level synthesis (SLS) parameterization, we propose a novel SLS design using an output-feedback affine control policy and extend it to a distributionally robust setting to improve system resilience by minimizing the cost function while ensuring constraint satisfaction against the worst-case uncertainty distribution. The scopes of model mismatch and stochastic disturbances are quantified using the 1-norm and a Wasserstein metric-based ambiguity set, respectively. For the closed-loop dynamics, we analyze the distributional shift between the predicted output-input response – computed using nominal parameters and empirical disturbance samples – and the actual closed-loop distribution, highlighting its dependence on model mismatch and SLS parameterization. Assuming convex and Lipschitz continuous cost functions and constraints, we derive a tractable reformulation of the distributionally robust SLS (DR-SLS) problem by leveraging tools from robust control and distributionally robust optimization (DRO). Numerical experiments validate the performance and robustness of the proposed approach.
@@ 1 Introduction
To enhance robustness and reduce conservatism in optimal control design, closed-loop feedback policies are generally preferred over open-loop control actions in both robust optimal decision-making and Model Predictive Control (MPC) frameworks [4, 17]. However, incorporating feedback policies directly into online MPC optimization introduces significant computational challenges, as it results in nonconvex optimization problems–even when the system dynamics and control laws are linear. For linear systems, the system level synthesis (SLS) framework provides a parameterization of the closed-loop system. Other than the conventional MPC design focusing on optimizing controllers, SLS-based design alternatively focuses on directly optimizing the whole closed-loop performance by solving convex optimization problems and reconstructing control policies from the optimized closed-loop response via linear transformations [18]. This shift in perspective has facilitated progress in distributed and robust predictive control design [18, 19, 20].
@@
Motivated by the above limitations, this paper investigates SLS design with output feedback affine control policies for uncertain linear systems subject to model mismatch and additive disturbances. This setting presents several major challenges: 1) the use of affine output feedback policies complicates the construction of an SLS parameterization while maintaining computational tractability; 2) model mismatch and stochastic disturbances lead to inaccurate predictions of system dynamics, increasing the difficulty of designing effective feedback policies; 3) model mismatch induces distributional shifts between the true closed-loop behavior and the predicted closed-loop response, further complicating the solution of the DR control formulation.
In order to address the above challenges, we propose a novel DR-SLS method with output feedback affine policies. The contributions of this work are summarized as follows:
@@
  • A computationally tractable relaxation of the proposed DR-SLS problem is developed by incorporating tools from robust SLS and DRO, resulting in a convex reformulation. Simulation results are provided to validate the effectiveness of the proposed approach.
@@ 2 Problem Formulation
where $x_{t}\in\mathbb{R}^{n}$ is the system state, $u_{t}\in\mathbb{R}^{m}$ is the control input, $e_{t}\in\mathbb{R}^{q}$ is the innovation process, and subscript $t$ is time index. The matrices $A$, $B$, $L$, $C$ and $D$ are the respective system, input, Kalman, output and feedforward matrices, respectively. More details about the innovation form of linear systems can be found in [26, 27, 28].
@@
where $T$ is the length of prediction horizon, $h_{t}(\cdot)$ and $g_{t}(\cdot)$ are stage cost and constraint functions, respectively, $\pi(\cdot)$ is the control policy, which is restricted to be strictly causal. In our design, $h_{t}$ and $g_{t}$ are assumed to be convex and Lipschitz continuous functions.
@@
In the upcoming sections, the SLS design for system (1) under control law (3) will be investigated. It is important to note that incorporating the affine control policy introduces computational challenges for existing SLS methods. Specifically, the presence of the bias term in the affine policy renders a direct extension of the existing SLS methods developed for linear control policies nonconvex, see [22, 23], and thus computationally demanding. In addition to developing an SLS design for system (1) under the assumption of perfect knowledge of system models, we further investigate the impact of model uncertainties for the proposed SLS design. To address the influence of model uncertainties, we propose a novel DR-SLS framework to improve the robustness of the optimal control solution against model mismatch and stochastic disturbances.
@@ 3.1 System Level Synthesis With Exact Model
and for any $\tau\in\mathbb{Z}_{+}$ (also called past horizon), we see that
@@
Supposing that (ALC)(A-LC) is Schur stable and denoting $\lambda_{\max}$ ($|\lambda_{\max}|<1$) as the eigenvalue of (ALC)(A-LC) with the largest absolute value, it was shown in [26] that
@@
$\mathbf{G}=$ $\mathbf{CT}_{u}+\mathbf{D},$ (10a)
$\mathbf{y}_{0}=$ $\mathbf{C}\Gamma\Psi\begin{bmatrix}\mathbf{u}_{0}^{-}\\ \mathbf{y}_{0}^{-}\end{bmatrix}$, (10b)
$\Theta=$ $\mathbf{CT}_{e}+I.$ (10c)
@@
The proof of the two equalities in (18a) follows a standard SLS design with linear policies, as shown in [25, 22]. By substituting $\mathbf{K}=\Phi_{u}\Phi_{y}^{-1}$, we obtain
@@
where the second equality exploits the property $\mathbf{G}\Phi_{u}=\Phi_{y}-I$ in (14). Similarly, one has
@@
Unlike in the existing SLS design with linear policies, the parameters $\phi_{y}$ and $\phi_{u}$ in (18b) quantify the contribution of the bias term of the affine policy to the closed-loop output and input responses, respectively. Substituting $\mathbf{p}=\phi_{u}-\mathbf{K}\phi_{y}$ into the first equality in (18b) gives
$\hskip-5.0pt(I-\mathbf{GK})^{-1}\mathbf{Gp}=$ $(I-\mathbf{GK})^{-1}\mathbf{G}(\phi_{u}-\mathbf{K}\phi_{y})$ (21a)
$=$ $(I-\mathbf{GK})^{-1}(\phi_{y}-\mathbf{GK}\phi_{y})=\phi_{y},$ (21b)
where the second equality builds upon the property $\mathbf{G}\phi_{u}=\phi_{y}$ in (14).
@@
$(I-\mathbf{KG})^{-1}\mathbf{p}=$ $(I+\mathbf{K}(I-\mathbf{GK})^{-1}\mathbf{G})\mathbf{p}$
$=$ $\mathbf{p}+\mathbf{K}(I-\mathbf{GK})^{-1}\mathbf{Gp}$
$=$ $\mathbf{p}+\mathbf{K}\phi_{y}=\phi_{u},$
where the first equality exploits the matrix inverse lemma (Sherman-Morrison-Woodbury formula) [30], the second equality adopts the results derived in (21), and the last equality utilizes the relationship in (17). This completes the proof. $\square$
@@
Remark 3: The existing SLS design approach with linear policies, e.g., [25], could be extended to affine policies by augmenting the output signal and feedback matrix as $\bar{y}_{t}:=[y_{t}^{\mathrm{T}},1]^{\mathrm{T}}$ and $\bar{K}_{t,k}=[K_{t,k},p_{k}^{\mathrm{T}}]$, respectively. However, this extension introduces potential issues. Specifically, considering the augmented output signal and feedback gains causes the matrix $\mathbf{K}$ in (12) to be block-lower-triangular rather than strictly block-lower-triangular. Consequently, the invertibility of $(I-\mathbf{GK})$, which is required in (18), is no longer guaranteed. However, if the system (1) is strictly proper, i.e., $D=0$, the matrix $\mathbf{G}$ defined in (10) becomes strictly-block-lower-triangular, and the existing approach with linear control policy is applicable to the extended system.
@@
where $\mathbb{P}^{\mathcal{M}}$ denotes the probability distribution of the closed-loop output-input response $\bm{\eta}$ with the system model $\mathcal{M}$, which influences the parameters $(\mathbf{G},\Theta,\mathbf{y}_{0},\mathbf{e})$, the functions $h(\cdot)$ and $g(\cdot)$ are constructed based on $h_{t}(\cdot)$ and $g_{t}(\cdot)$ defined in (2), respectively. The stochastic uncertainty considered in the above formulation is the innovation process vector $\mathbf{e}$.
@@ 3.2 SLS Parameterization with Model Mismatch
The SLS formulation proposed in Section 3.A is based on the assumption that the system model, i.e., the set of parameters $(\mathbf{G},\Theta,\mathbf{y}_{0},\mathbf{e})$, is perfectly known. However, obtaining the exact model information is nontrivial and challenging. It is more practical to consider that only an approximate model is available for SLS design.
It is worth noting that among the above-listed parameters, $\mathbf{G}$ plays the most important role in influencing the SLS design since it will directly influence the SLS design via (14). In this section, the SLS formulation proposed in Theorem 1 of Section 3.A is extended to the case where only the approximated parameter $\hat{\mathbf{G}}$ is available. For ease of notation, the approximation error of $\mathbf{G}$ is defined as $\Delta:=\mathbf{G}-\hat{\mathbf{G}}$.
@@
Proof: Based on Theorem 1, we have
$\begin{bmatrix}I&-\mathbf{\hat{G}}\end{bmatrix}\begin{bmatrix}\hat{\Phi}_{y}\\ \hat{\Phi}_{u}\end{bmatrix}=I,$
@@
$\begin{bmatrix}I&-\mathbf{G}\end{bmatrix}\begin{bmatrix}\hat{\Phi}_{y}\\ \hat{\Phi}_{u}\end{bmatrix}=I-\Delta\hat{\Phi}_{u}.$
Consequently, we arrive at
$\begin{bmatrix}I&-\mathbf{G}\end{bmatrix}\begin{bmatrix}\hat{\Phi}_{y}\\ \hat{\Phi}_{u}\end{bmatrix}\left(I-\Delta\hat{\Phi}_{u}\right)^{-1}=I,$ (26)
where the invertibility of $(I-\Delta\hat{\Phi}_{u})$ is guaranteed since $\Delta$ is block-lower-triangular and $\hat{\Phi}_{u}$ is strictly-block-lower-triangular. The above equality provides an option for selecting $\Phi_{y}=\hat{\Phi}_{y}(I-\Delta\hat{\Phi}_{u})^{-1}$ and $\Phi_{u}=\hat{\Phi}_{u}(I-\Delta\hat{\Phi}_{u})^{-1}$ that ensures $\mathbf{K}=\Phi_{u}\Phi_{y}^{-1}=\hat{\Phi}_{u}\hat{\Phi}_{y}^{-1}$.
Since $\hat{\phi}_{y}$ and $\hat{\phi}_{u}$ satisfy
@@
where the relationship in (29b) is built upon the fact that the control policy parameter $\mathbf{p}$ is identical in both cases of the true closed-loop parameterization and nominal closed-loop parameterization.
Denoting $r_{y}:=\phi_{y}-\hat{\phi}_{y}$ and $r_{u}:=\phi_{u}-\hat{\phi}_{u}$, and combining (28) and (29) results in
@@ 4.1 Distributionally Robust SLS Formulation
Solving the SLS-based problem (23) is generally intractable when assuming the existence of model mismatch for parameters $(\mathbf{G},\Theta,\mathbf{y}_{0})$, and uncertain probability distribution $\mathbb{P}_{\mathbf{e}}$. One possible solution for solving (23) is adopting the sample average approximation (SAA) and certainty equivalence (CE) principles, in which the unknown system parameters $(\mathbf{G},\Theta,\mathbf{y}_{0})$ are replaced by their nominal values $(\hat{\mathbf{G}},\hat{\Theta},\hat{\mathbf{y}}_{0})$ and the true innovation process distribution $\mathbb{P}_{\mathbf{e}}$ is approximated by its empirical counterpart $\overline{\mathbb{P}}_{\mathbf{e}}$. In this way, the true closed-loop response distribution $\mathbb{P}^{\mathcal{M}}$ is replaced by its empirical approximated distribution $\overline{\mathbb{P}}^{\hat{\mathcal{M}}}$.
@@
In the following, a distributionally robust SLS (DR-SLS) design will be presented, considering the uncertainties introduced by the model mismatch between $(\mathbf{G},\Theta,\mathbf{y}_{0})$ and $(\hat{\mathbf{G}},\hat{\Theta},\hat{\mathbf{y}}_{0})$ as well as the estimation error of the empirical innovation process distribution $\overline{\mathbb{P}}_{\mathbf{e}}$. The approximation errors of $(\mathbf{G},\Theta,\mathbf{y}_{0})$ are denoted as $\Delta:=\mathbf{G}-\hat{\mathbf{G}}$, $\tilde{\Theta}:=\Theta-\hat{\Theta}$, and $\tilde{\mathbf{y}}_{0}:=\mathbf{y}_{0}-\hat{\mathbf{y}}_{0}$, respectively. Before presenting the main results of this section, the following definitions are introduced.
@@
Given $N$ samples of the innovation process $\{\mathbf{e}_{i}\}_{i=1}^{N}$, the empirical distribution $\overline{\mathbb{P}}_{\mathbf{e}}$ can be constructed as $\overline{\mathbb{P}}_{\mathbf{e}}:=\frac{1}{N}\sum_{i=1}^{N}\delta(\mathbf{e}_{i})$, where $\delta(\cdot)$ is the Dirac distribution. Accordingly, the empirical predictive distribution of the output-input response $\bm{\eta}$ with the nominal model $\mathcal$$\hat${M}} is